# Adapting Newton’s Method to Neural Networks through a Summary of Higher-Order Derivatives

### v3

@@ 4.2 Regularizing $\bar{\mathbf{H}}$ by using order-3 information
By using the technique presented in Section 3, the diagonal coefficients $(D_{1},\cdots,D_{S})$ of $\mathbf{D}_{\bm{\theta}}^{(3)}(\mathbf{u})\in\mathbb{R}^{S\times S\times S}$ are available with little computational cost. Let: $\mathbf{D}:=\mathrm{Diag}(|D_{1}|^{1/3},\cdots,|D_{S}|^{1/3})\in\mathbb{R}^{S}$.
We modify the method of Nesterov & Polyak (2006) by integrating an anisotropic factor $\mathbf{D}$ into the cubic term. Thus, our goal is to minimize according to $\bm{\eta}$ the function $T$: $T(\bm{\eta}):=-\bm{\eta}^{T}\bar{\mathbf{g}}+\frac{1}{2}\bm{\eta}\bar{\mathbf{H}}\bm{\eta}+\frac{\lambda_{\mathrm{int}}}{6}\|\mathbf{D}\bm{\eta}\|^{3}$, where $\lambda_{\mathrm{int}}$ is the internal damping coefficient, which can be used to tune the strength of the cubic regularization. Under conditions detailed in Appendix D, this minimization problem is equivalent to finding a solution $\bm{\eta}_{*}$ such that:
$\bm{\eta}_{*}=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\bm{\eta}_{*}\|\mathbf{D}^{2}\right)^{-1}\bar{\mathbf{g}},$ (10)
which is no more than a regularized version of 8. Finally, this multi-dimensional minimization problem boils down to a scalar root finding problem (see Appendix D).
@@ 4.3 Properties
The final method is a combination of the training step 8 with regularization 10:
Method 1.
Training step $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}^{*}$, where $\bm{\eta}_{t}^{*}$ is the solution with the largest norm $\|\mathbf{D}_{t}\bm{\eta}\|$ of the equation: $\bm{\eta}=\left(\bar{\mathbf{H}}_{t}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}_{t}\bm{\eta}\|\mathbf{D}_{t}^{2}\right)^{-1}\bar{\mathbf{g}}_{t}$.
@@
Without the cubic regularization ($\lambda_{\mathrm{int}}=0$), Newton’s method is recovered when using the discrete partition, that is, $S=P$ with $\mathcal{I}_{s}=\{s\}$ for all $s$, and Cauchy’s steepest descent is recovered when using the trivial partition, that is, $S=1$ with $\mathcal{I}_{1}=\{1,\cdots,P\}$. See Appendix C for more details.
@@
Using Equations 8 or 10 requires solving only a linear system of $S$ equations, instead of $P$ in Newton’s method. With the cubic regularization, only a constant term is added to the complexity, since it is a matter of scalar root finding.
@@
As showed in Appendix E, under a condition on the directions $\mathbf{u}_{t}$, [^4] This holds typically if $\mathbf{u}_{t}$ is the gradient or a moving average of the gradients (momentum). the trajectory of optimization of a model trained by Method 1 is invariant by affine reparameterization of the sub-vectors of parameters $\bm{\theta}_{\mathcal{I}_{s}}:=\mathrm{vec}(\{\theta_{p}:p\in\mathcal{I}_{s}\})$. Let $(\alpha_{s})_{1\leq s\leq S}$ and $(\beta_{s})_{1\leq s\leq S}$ be a sequence of nonzero scalings and a sequence of offsets, and $\tilde{\bm{\theta}}$ such that, for all $1\leq s\leq S$, $\tilde{\bm{\theta}}_{\mathcal{I}_{s}}=\alpha_{s}\bm{\theta}_{\mathcal{I}_{s}}+\beta_{s}$. Then, the training trajectory of the model is the same with both parameterizations $\bm{\theta}$ and $\tilde{\bm{\theta}}$. This property is desirable in the case of neural networks, where one can use either the usual or the NTK parameterization, which consists of a layer-wise scaling of the parameters. The relevance of this property is discussed in Appendix E.1.
Compared to the standard regularization $\bar{\mathbf{H}}+\lambda\mathbf{I}$ and Nesterov’s cubic regularization, the anisotropic Nesterov regularization does not break the property of invariance by subset-wise scaling of the parameters of 8. This is mainly due to our choice to keep only the diagonal coefficients of $\mathbf{D}_{\bm{\theta}}^{(3)}(\mathbf{u})$ while discarding the others. In particular, the off-diagonal coefficients contain cross-derivatives that would be difficult to include in an invariant training step.
@@ 5 Experiments
5.1 Empirical computation of $\bar{\mathbf{H}}$ and $\bm{\eta}$
As recalled in Section 2, many works perform a diagonal, block-diagonal or block-tridiagional (Martens & Grosse, 2015) approximation of the Hessian or its inverse. Since a summary $\bar{\mathbf{H}}$ of the Hessian and its inverse $\bar{\mathbf{H}}^{-1}$ are available and all their off-diagonal coefficients have been computed and kept, one can to check if these coefficients are indeed negligible.
@@ 5.1 Empirical computation of $\bar{\mathbf{H}}$ and $\bm{\eta}$
We have trained LeNet-5 and VGG-11’ [^5] VGG-11’ is a variant of VGG-11 with only one fully-connected layer at the end, instead of 3. on CIFAR-10 using SGD with momentum. Before each epoch, we compute the full-batch gradient, denoted by $\mathbf{u}$, which we use as a direction to compute $\bar{\mathbf{H}}$, again in full-batch. We report submatrices of $\bar{\mathbf{H}}$ and $\bar{\mathbf{H}}^{-1}$ at initialization and at the epoch where the validation loss is the best in Figure 1(a) (LeNet) and Figure 1(b) (VGG-11’).
@@
First, the block-diagonal approximation of the Hessian is indeed very rough, while the block-diagonal approximation of the inverse Hessian seems to be more reasonable (at least in these setups), which has already been shown by Martens & Grosse (2015). Second, there seem to be long-range interactions between layers, both at initialization and after several epochs. For LeNet, all the layers (except the first one) seem to interact together at initialization (Fig. 1(a)). In the matrix $\bar{\mathbf{H}}^{-1}$ computed on VGG, the last 3 layers interact strongly and the last 6 layers also interact, but a bit less.
@@
Figure 1: Setup: models trained by SGD on CIFAR-10. Submatrices of $\bar{\mathbf{H}}$ (first row) and $\bar{\mathbf{H}}^{-1}$ (second row), where focus is on interactions: weight-weight, weight-bias, bias-bias of the different layers, at initialization and before best validation loss epoch.
Results on $\bm{\eta}_{*}$.
The evolution of the learning rates $\eta_{*}$ computed according to 10 in LeNet and VGG is shown in Figure 2(b). First, the learning rates computed for the biases are larger than those computed for the weights. Second, even if only the weights are considered, the computed $\eta_{*}$ can differ by several orders of magnitude. Finally, the first two layers of LeNet (which are convolutional) have smaller $\eta_{*}$ than the last three layers (which are fully-connected). Conversely, in VGG, the weights of the last (convolutional) layers have a smaller $\eta_{*}$ than those of the first layers.
@@ 5.2 Training experiments
In this section, we show a proof-of-concept of the optimization method 1 (summarized in Algorithm 1), on simple vision tasks and with medium-sized neural networks. All the implementation details are available in Appendix G. In particular, we have introduced a step size $\lambda_{1}$ that leads to the following modification of the training step 8: $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\lambda_{1}\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}^{*}$.
Algorithm 1 Informal implementation of the second-order method described in Section 4.
@@
     $\mathbf{g}_{t}\leftarrow\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta}_{t},Z_{t})$ (backward pass)
@@
     $\bar{\mathbf{g}}_{t}\leftarrow\mathbf{D}_{\bm{\theta}_{t}}^{(1)}(\mathbf{u}_{t})=\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta}_{t},\tilde{Z}_{t})$
     $\bar{\mathbf{H}}_{t}\leftarrow\mathbf{D}_{\bm{\theta}_{t}}^{(2)}(\mathbf{u}_{t})=\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\bm{\theta}^{2}}(\bm{\theta}_{t},\tilde{Z}_{t})\mathbf{U}_{t}\mathbf{I}_{P:S}$
     $\mathbf{D}_{t}\leftarrow\mathrm{Diag}(|\mathbf{D}_{\bm{\theta}_{t}}^{(3)}(\mathbf{u}_{t})|_{iii}^{1/3}:i\in\{1,\cdots,S\})\in\mathbb{R}^{S^{2}}$
     $\bm{\eta}_{t}\leftarrow$ solution of $\bm{\eta}=\left(\bar{\mathbf{H}}_{t}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}_{t}\bm{\eta}\|\mathbf{D}_{t}^{2}\right)^{-1}\bar{\mathbf{g}}_{t}$ with max. norm $\|\mathbf{D}_{t}\bm{\eta}\|$ (Method 1)
     $\bm{\theta}_{t+1}\leftarrow\bm{\theta}_{t}-\lambda\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}_{t}$ (training step)
@@
  • LeNet: LeNet-5 (LeCun et al., 1998) model trained on CIFAR-10 with 2 convolutional layers of sizes 6, 16, and 3 fully connected layers of sizes 120, 84, 10;
  • VGG: VGG-11’ trained on CIFAR-10. VGG-11’ is a variant of VGG-11 (Simonyan & Zisserman, 2014) with only one fully-connected layer at the end, instead of 3, with $\mathrm{ELU}$ activation function (Clevert et al., 2015), without batch-norm;
@@
  • Adam: the best learning rate has been selected by grid-search;
  • K-FAC: the best learning rate and damping have been selected by grid-search;
  • NewtonSummary (ours): the best $\lambda_{1}$ and $\lambda_{\mathrm{int}}$ have been selected by grid search.
@@ Results.
The evolution of the training loss is plotted in Figure 2(a) for each of the 3 optimization methods, for 5 different seeds. In each set of experiments, the training is successful, but slow or unstable at some points. Anyway, the minimum training loss achieved by Method 1 (NewtonSummary) is comparable to the minimum training loss achieved by K-KAC or Adam in all the series except for BigMLP, whose training is extremely slow. We provide the results on the test set in Appendix I and a comparison of the training times in Appendix J.
Some runs have encountered instabilities due to very large step sizes $\bm{\eta}_{*}$. In fact, we did not use any safeguards, such as a regularization term $\lambda\mathbf{I}$ added to $\bar{\mathbf{H}}$, or clipping the learning rates to avoid increasing the number of hyperparameters.
@@
(b) Setup: LeNet, VGG-11’ trained by SGD on CIFAR-10. Learning rates $\eta_{*}$ computed according to 10, specific to each tensor of weights and tensor of biases of each layer. For each epoch $k\in\{10,30,50,70,90\}$, the reported value has been averaged over the epochs $[k-10,k+9]$ to remove the noise.
@@
Since the matrix $\bar{\mathbf{H}}$ can be computed numerically as long as $S$ remains relatively small, this method may become unpractical for very large models. However, Method 1 is flexible enough to be adapted to such models: one can regroup tensors “of the same kind” to build a coarser partition of the parameters, and thus obtain a small $S$, which is exactly what is needed to compute $\bar{\mathbf{H}}$ and invert it. The difficulty would then be to find a good partition of the parameters, by grouping all the tensors that “look alike”. We provide an example in Appendix H with a very deep multilayer perceptron.
@@
We propose in Appendix K an empirical study and a discussion about the choice of the partition of the parameters. We show how it affects the training time and the final loss.
@@
We show in Appendix L that the interactions between layers cannot be neglected when using our method: Method 1 outperforms its diagonal approximation on LeNet and VGG11’, showing the importance of off-diagonal coefficients of $\bar{\mathbf{H}}$.

### v4

@@ 5.1 Presentation
By using the technique presented in Section 3, the diagonal coefficients $(D_{1},\cdots,D_{S})$ of $\mathbf{D}_{\boldsymbol{\theta}}^{(3)}(\mathbf{u})\in\mathbb{R}^{S\times S\times S}$ are available with little computational cost. Let $\mathbf{D}:=\mathrm{Diag}(|D_{1}|^{1/3},\cdots,|D_{S}|^{1/3})\in\mathbb{R}^{S}$.
We modify the method of [24] by integrating an anisotropic factor $\mathbf{D}$ into the cubic term. Thus, our goal is to minimize according to $\boldsymbol{\eta}$ the function $T$: $T(\boldsymbol{\eta}):=-\boldsymbol{\eta}^{T}\bar{\mathbf{g}}+\frac{1}{2}\boldsymbol{\eta}\bar{\mathbf{H}}\boldsymbol{\eta}+\frac{\lambda_{\mathrm{int}}}{6}\|\mathbf{D}\boldsymbol{\eta}\|^{3}$, where $\lambda_{\mathrm{int}}$ is the internal damping coefficient, which can be used to tune the strength of the cubic regularization. Under conditions detailed in Appendix 0.D, this minimization problem is equivalent to finding a solution $\boldsymbol{\eta}_{*}$ such that:
$\boldsymbol{\eta}_{*}=\left(\bar{\mathbf{H}}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}\boldsymbol{\eta}_{*}\|\mathbf{D}^{2}\right)^{-1}\bar{\mathbf{g}},$ (10)
which is a regularized version of (8). Finally, this multi-dimensional minimization problem boils down to a scalar root finding problem (see Appendix 0.D).
@@ 5.2 Properties
The final method is a combination of the learning rate computed in Eqn. (8) with regularization (10):
Method 1
Training step $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}^{*}$, where $\boldsymbol{\eta}_{t}^{*}$ is the solution with the largest norm $\|\mathbf{D}_{t}\boldsymbol{\eta}\|$ of the equation: $\boldsymbol{\eta}=\left(\bar{\mathbf{H}}_{t}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}_{t}\boldsymbol{\eta}\|\mathbf{D}_{t}^{2}\right)^{-1}\bar{\mathbf{g}}_{t}$.
@@
Without the cubic regularization ($\lambda_{\mathrm{int}}=0$), Newton’s method is recovered when using the discrete partition, that is, $S=P$ with $\mathcal{I}_{s}=\{s\}$ for all $s$, and Cauchy’s steepest descent is recovered when using the trivial partition, that is, $S=1$ with $\mathcal{I}_{1}=\{1,\cdots,P\}$. See Appendix 0.C for more details.
@@
Using Equations (8) or (10) requires solving only a linear system of $S$ equations, instead of $P$ in Newton’s method. With the cubic regularization, only a constant term is added to the complexity, since it is a matter of scalar root finding.
@@
As showed in Appendix 0.E, under a condition on the directions $\mathbf{u}_{t}$, [^4] It holds if $\mathbf{u}_{t}$ is the gradient or a moving average of the gradients (momentum). the trajectory of optimization of a model trained by Method 1 is invariant by affine reparameterization of the sub-vectors of parameters $\boldsymbol{\theta}_{\mathcal{I}_{s}}:=\mathrm{vec}(\{\theta_{p}:p\in\mathcal{I}_{s}\})$. Let $(\alpha_{s})_{1\leq s\leq S}$ and $(\beta_{s})_{1\leq s\leq S}$ be a sequence of nonzero scalings and a sequence of offsets, and $\tilde{\boldsymbol{\theta}}$ such that, for all $1\leq s\leq S$, $\tilde{\boldsymbol{\theta}}_{\mathcal{I}_{s}}=\alpha_{s}\boldsymbol{\theta}_{\mathcal{I}_{s}}+\beta_{s}$. Then, the training trajectory of the model is the same with both parameterizations $\boldsymbol{\theta}$ and $\tilde{\boldsymbol{\theta}}$. This property is desirable in the case of neural networks, where one can use either the usual or the NTK parameterization, which consists of a layer-wise scaling of the parameters. The relevance of this property is discussed in Appendix 0.E.1.
Compared to the standard regularization $\bar{\mathbf{H}}+\lambda\mathbf{I}$ and Nesterov’s cubic regularization, the anisotropic Nesterov regularization does not break the property of invariance by subset-wise scaling of the parameters of (8). This is mainly due to our choice to keep only the diagonal coefficients of $\mathbf{D}_{\boldsymbol{\theta}}^{(3)}(\mathbf{u})$ while discarding the others. In particular, the off-diagonal coefficients contain cross-derivatives that would be difficult to include in an invariant training step.
@@ 6 Experiments
6.1 Empirical computation of $\bar{\mathbf{H}}$ and $\boldsymbol{\eta}_{*}$
As recalled in Section 2, many works perform a diagonal, block-diagonal or block-tridiagional [20] approximation of the Hessian or its inverse. Since a summary $\bar{\mathbf{H}}$ of the Hessian and its inverse $\bar{\mathbf{H}}^{-1}$ are available and all their off-diagonal coefficients have been computed and kept, one can to check if these coefficients are indeed negligible.
@@ 6.1 Empirical computation of $\bar{\mathbf{H}}$ and $\boldsymbol{\eta}_{*}$
We have trained LeNet-5 and VGG-11’ [^5] VGG-11’ is a variant of VGG-11 with 1 final fully-connected layer instead of 3. on CIFAR-10 using SGD with momentum. Before each epoch, we compute the full-batch gradient, denoted by $\mathbf{u}$, which we use as a direction to compute $\bar{\mathbf{H}}$, again in full-batch. We report submatrices of $\bar{\mathbf{H}}$ and $\bar{\mathbf{H}}^{-1}$ at initialization and at the epoch where the validation loss is the best in Figure 1(a) (LeNet) and Figure 1(b) (VGG-11’).
@@
First, the block-diagonal approximation of the Hessian is indeed very rough, while the block-diagonal approximation of the inverse Hessian seems to be more reasonable (at least in these setups), which has already been shown by [20]. Second, there seem to be long-range interactions between layers, both at initialization and after several epochs. For LeNet, all the layers (except the first one) seem to interact together at initialization (Fig. 1(a)). In the matrix $\bar{\mathbf{H}}^{-1}$ computed on VGG, the last 3 layers interact strongly and the last 6 layers also interact, but a bit less.
@@
Figure 1: Setup: models trained by SGD on CIFAR-10. Submatrices of $\bar{\mathbf{H}}$ (1st row) and $\bar{\mathbf{H}}^{-1}$ (2nd row), where focus is on interactions: weight-weight, weight-bias, bias-bias of the different layers, at initialization and before best validation loss.
Results on $\boldsymbol{\eta}_{*}$.
The evolution of the learning rates $\eta_{*}$ computed according to (10) in LeNet and VGG is shown in Figure 2(b). First, the learning rates computed for the biases are larger than those computed for the weights. Second, even if only the weights are considered, the computed $\eta_{*}$ can differ by several orders of magnitude. Finally, the first two layers of LeNet (which are convolutional) have smaller $\eta_{*}$ than the last three layers (which are fully-connected). Conversely, in VGG, the weights of the last (convolutional) layers have a smaller $\eta_{*}$ than those of the first layers.
@@ 6.2 Training experiments
To show that the projections of the 2nd and 3rd order derivatives of the loss defined in Section 3 can be practically used to train neural networks, we test our optimization method 1 (summarized in Algorithm 1) on simple vision tasks. All the implementation details are available in Appendix 0.G. In particular, we have introduced a step size $\lambda_{1}$ that leads to the following modification of the training step (8): $\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\lambda_{1}\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}^{*}$.
Algorithm 1 Informal description of the 2nd-order method described in Sec. 5.
@@
     $\mathbf{g}_{t}\leftarrow\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta}_{t},Z_{t})$ (backward pass)
@@
     $\bar{\mathbf{g}}_{t}\leftarrow\mathbf{D}_{\boldsymbol{\theta}_{t}}^{(1)}(\mathbf{u}_{t})=\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta}_{t},\tilde{Z}_{t})$
     $\bar{\mathbf{H}}_{t}\leftarrow\mathbf{D}_{\boldsymbol{\theta}_{t}}^{(2)}(\mathbf{u}_{t})=\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{2}}(\boldsymbol{\theta}_{t},\tilde{Z}_{t})\mathbf{U}_{t}\mathbf{I}_{P:S}$
     $\mathbf{D}_{t}\leftarrow\mathrm{Diag}(|\mathbf{D}_{\boldsymbol{\theta}_{t}}^{(3)}(\mathbf{u}_{t})|_{iii}^{1/3}:i\in\{1,\cdots,S\})\in\mathbb{R}^{S^{2}}$
     $\boldsymbol{\eta}_{t}\leftarrow$ sol. of $\boldsymbol{\eta}=\left(\bar{\mathbf{H}}_{t}+\frac{\lambda_{\mathrm{int}}}{2}\|\mathbf{D}_{t}\boldsymbol{\eta}\|\mathbf{D}_{t}^{2}\right)^{-1}\bar{\mathbf{g}}_{t}$ with max. norm $\|\mathbf{D}_{t}\boldsymbol{\eta}\|$ (Method 1)
     $\boldsymbol{\theta}_{t+1}\leftarrow\boldsymbol{\theta}_{t}-\lambda\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}_{t}$ (training step)
@@
  • LeNet: LeNet-5 [16] model trained on CIFAR-10 with 2 convolutional layers of sizes 6, 16, and 3 fully connected layers of sizes 120, 84, 10;
  • VGG: VGG-11’ trained on CIFAR-10. VGG-11’ is a variant of VGG-11 [31] with only one fully-connected layer at the end, instead of 3, with $\mathrm{ELU}$ activation function [5], without batch-norm;
@@
  • Adam: learning rate selected by grid-search;
  • K-FAC: learning rate and damping selected by grid-search;
  • NewtonSummary (ours): $\lambda_{1}$ and $\lambda_{\mathrm{int}}$ selected by grid search.
@@ Results.
The evolution of the training loss is plotted in Figure 2(a) for each of the 3 optimization methods, for 5 different seeds. In each set of experiments, the training is successful, but slow or unstable at some points. Anyway, the minimum training loss achieved by Method 1 (NewtonSummary) is comparable to the minimum training loss achieved by K-KAC or Adam in all the series except for MLP, whose training is slower. We provide the results on the test set in Appendix 0.I and a comparison of the training times in Appendix 0.M.
Some runs have encountered instabilities due to very large step sizes $\boldsymbol{\eta}_{*}$. In fact, we did not use any safeguards, such as a regularization term $\lambda\mathbf{I}$ added to $\bar{\mathbf{H}}$, or clipping the learning rates to avoid increasing the number of hyperparameters.
@@
(b) Setup: LeNet, VGG-11’ trained by SGD on CIFAR-10. Learning rates $\eta_{*}$ computed according to (10), specific to each tensor of weights and tensor of biases of each layer. For each epoch $k\in\{10,30,50,70,90\}$, the reported value has been averaged over the epochs $[k-10,k+9]$ to remove the noise.
@@
Since the matrix $\bar{\mathbf{H}}$ can be computed numerically as long as $S$ remains relatively small, this method may become unpractical for very large models. However, Method 1 is flexible enough to be adapted to such models: one can regroup tensors “of the same kind” to build a coarser partition of the parameters, and thus obtain a small $S$, which is exactly what is needed to compute $\bar{\mathbf{H}}$ and invert it. The difficulty would then be to find a good partition of the parameters, by grouping all the tensors that “look alike”. We provide an example in Appendix 0.H with a very deep multilayer perceptron.
@@
We propose in Appendix 0.J an empirical study and a discussion about the choice of the partition of the parameters. We show how it affects the training time and the final loss.
@@
We show in Appendix 0.K that the interactions between layers cannot be neglected when using our method: Method 1 outperforms its diagonal approximation on LeNet and VGG11’, showing the importance of off-diagonal coefficients of $\bar{\mathbf{H}}$.